OpenStudy (anonymous):

RST is dilated about the origin, O, to create ΔR'S'T'. Point R is located at (3, 9), and point R' is located at (1.2, 3.6). Which scale factor was used to perform the dilation?

OpenStudy (anonymous):

Are there any answer choices?

OpenStudy (anonymous):

no I have to fill it in myself...

OpenStudy (anonymous):

To find the scale factor, divide the first coordinates by the second coordinates: x-coordinates: 3÷1.2=2.5 y-coordinates: 9÷3.6=2.5 The scale factor is: ? @shortyyyy99

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

No, it's not 0.4

OpenStudy (anonymous):

It's 2.5

OpenStudy (anonymous):

it is 0.4... that's the answer I was given...

OpenStudy (anonymous):

By who?

OpenStudy (anonymous):

To find the scale factor, divide the first coordinates by the second coordinates: x-coordinates: 3÷1.2 =2.5 y-coordinates: 9÷3.6=2.5 The scale factor is: 2.5

OpenStudy (anonymous):

by the lesson I was taking .